OpenStudy (anonymous):

The width of a rectangle is 6 in. less than its length. The perimeter is 68 in. What is the width of the rectangle? ____?___ in.

OpenStudy (anonymous):

@Meehan98 plz help me

OpenStudy (anonymous):

@JB1625 plz help me

OpenStudy (anonymous):

this is the hardest question on the test

OpenStudy (anonymous):

148?

OpenStudy (meehan98):

So if you're width = 20, that would satisfy the equation because perimeter is all the sides added together. (20+20+14+14=68)

OpenStudy (anonymous):

this is the way i did it: (68+68)+(6+6)=148. is this right?

OpenStudy (jb1625):

@TwilightSparkle2015 Ok now I can help

OpenStudy (meehan98):

Ugh..I messed up on the drawing.

OpenStudy (anonymous):

68+68+6+6=148

OpenStudy (meehan98):

Okay, did you get it figured out?